the lengths of the sides of a rectangle are whole numbers. if the rectangles perimeter is 24 units, which of the following could not be the area of the rectangle?
a. 27 square units
b. 24 square units
c. 20 square units
d. 11 square units

Answer by hamsanash1981@gmail.com(151)

Given Perimeter = 24 units
2(l+b)=24
L+B =12
Option A can be area of the rectangle as 9*3 is 27 and 9+3 = 12.
Option B cannot be area of the rectangle as possible combination area (8*3,12*2,6*4) and their sum is (11,14,10).
option C can be area as 10*2 is 20 and 10+2 = 12
option D can be area as 11*1 is 11 and 11+1 = 12